The Creative Publishing Company (CPC) is a coupon book publisher with markets in several
southeastern states. CPC coupon books are either sold directly to the public, sold through
religious and other charitable organizations, or given away as promotional items. Operating
experience during the past year suggests the following demand function for CPC’s coupon
books:
Q = 5,000 – 4,000P + 0.02Pop + 0.5I + 1.5A
Where: Q is quantity, P is price ($), Pop is population, I is disposable income per household ($),
and A is advertising expenditures ($).
A. Determine the demand faced by CPC in a typical market in which P = $10, Pop = 1,000,000
persons, I = $30,000, and A = $10,000.
B. Calculate the level of demand if CPC increases annual advertising expenditures from
$10,000 to $15,000.
C. Calculate the demand curves faced by CPC in parts A and B.
1
Expert's answer
2017-10-13T15:01:06-0400
Q = 5,000 – 4,000P + 0.02Pop + 0.5I + 1.5A
Where: Q is quantity, P is price ($), Pop is population, I is disposable income per household ($), and A is advertising expenditures ($).
A. The demand faced by CPC in a typical market in which P = $10, Pop = 1,000,000
persons, I = $30,000, and A = $10,000 is:
Qd = 5,000 – 4,000*10 + 0.02*1,000,000 + 0.5*30,000 + 1.5*10,000 = 15,000 units.
B. The level of demand if CPC increases annual advertising expenditures from $10,000 to $15,000 is:
Qd = 5,000 – 4,000*10 + 0.02*1,000,000 + 0.5*30,000 + 1.5*15,000 = 22,500 units.
C. The demand curves faced by CPC in parts A and B are:
Qd = 55,000 - 4,000P and Qd = 62,500 - 4,000P.